AnimeGhoul8863:

Let f(x) = 2x + 2. Solve f^−1(x) when x = 4.

AnimeGhoul8863:

@Vocaloid

NarutoandSasuke:

that's easy e.e

AnimeGhoul8863:

T~T

NarutoandSasuke:

what grade r u in anime

AnimeGhoul8863:

10th

NarutoandSasuke:

oo thats why

AnimeGhoul8863:

ok so this is what i gots so far

AnimeGhoul8863:

\[f(x) = 2x + 2\] \[f(4)=2(4) + 2\] \[y=2(4)+2\] \[4=2y+2 \] \[-2 -2\] \[\frac{4-2 }{ 2}=y\] \[\frac{ 4-2 }{ 2 }=f^-1(4)\] \[\frac{ 2 }{ 2 }=f^-1(4) \]

AnimeGhoul8863:

^what i have so far now im stuck

AnimeGhoul8863:

@TheSmartOne

TheSmartOne:

What kind of math are you even doing Oh my

AnimeGhoul8863:

Solving Inverses of Functions i think

AnimeGhoul8863:

tbh i dont know its my first day of algebra and i am confused to my max

TheSmartOne:

Wait I see what you were doing Yeah your answer is correct, but next time first find the inverse without plugging in the number because otherwise you can get confused and make mistakes You can simply 2/2 to 1

AnimeGhoul8863:

so the answer is 1ok makes a little more since can u check some more of mine